# An insight on some properties of high order nonstandard linear multistep methods

### v1

@@ 3.3 The choice of function $\varphi$
In Section 5 we perform several numerical experiments comparing these choices. We should also note here that in [49], Dang and Hoang demonstrated a different construction of functions fulfilling these conditions up to an arbitrary order.
@@ 4 Proofs of the statements of Section 3
We prove the two directions of the statement separately.
@@
    Let us start with the (C1$\Rightarrow$(C2) direction.
    We prove the statement by induction: first, we show that it holds for $n=0$ and $n=1$ (the case of $n=2$ is also presented for easier understanding), and then by assuming that it holds for $n=N$, we validate the case $n=N+1$.
    If $n=0$, (8) has the form $f(x)=f(a)+e_{0}(x)$, meaning that $\lim_{x\rightarrow a}e_{0}(x)=\lim_{x\rightarrow a}f(x)-f(a)=0$.
    If $n=1$, (8) has the form $f(x)=f(a)+f^{\prime}(a)\varphi(x-a)+e_{1}(x)\varphi(x-a)$, from which we get
    $\lim_{x\rightarrow a}e_{1}(x)=\lim_{x\rightarrow a}\dfrac{f(x)-\left(f(a)+f^{\prime}(a)\varphi(x-a)\right)}{\varphi(x-a)}=$
    Now we can use the rule of l’Hospital, since $\varphi(0)=0$:
    $=\lim_{x\rightarrow a}\dfrac{f^{\prime}(x)-f^{\prime}(a)\varphi^{\prime}(x-a)}{\varphi^{\prime}(x-a)}=0,$
    where we used that $\varphi^{\prime}(0)=1$.
    If $n=2$, (8) has the form
    $f(x)=f(a)+f^{\prime}(a)\varphi(x-a)+\dfrac{f^{\prime\prime}(a)}{2}\left(\varphi(x-a)\right)^{2}+e_{2}(x)\left(\varphi(x-a)\right)^{2},$
    from which we get
    $\lim_{x\rightarrow a}e_{2}(x)=\lim_{x\rightarrow a}\dfrac{f(x)-\left(f(a)+f^{\prime}(a)\varphi(x-a)+\frac{f^{\prime\prime}(a)}{2}\left(\varphi(x-a)\right)^{2}\right)}{\left(\varphi(x-a)\right)^{2}}=$
    Now we can use the rule of l’Hospital, since $\varphi(0)=0$:
    $=\lim_{x\rightarrow a}\dfrac{f^{\prime}(x)-f^{\prime}(a)\varphi^{\prime}(x-a)-f^{\prime\prime}(a)\varphi(x-a)\varphi^{\prime}(x-a)}{2\varphi(x-a)\varphi^{\prime}(x-a)}=$
    Here we use l’Hospital’s rule again:
    $=\lim_{x\rightarrow a}\dfrac{f^{\prime\prime}(x)-f^{\prime}(a)\varphi^{\prime\prime}(x-a)-f^{\prime\prime}(a)\left(\left(\varphi^{\prime}(x-a)\right)^{2}+\varphi(x-a)\varphi^{\prime\prime}(x-a)\right)}{2\left(\varphi^{\prime}(x-a)\right)^{2}+2\varphi(x-a)\varphi^{\prime\prime}(x-a)}=0,$
    where we used that $\varphi(0)=0$, $\varphi^{\prime}(0)=1$ and $\varphi^{\prime\prime}(0)=0$.
    Now let us assume that (C2) holds for $n=N$. Observe that it is enough to show that
    $\lim_{x\rightarrow a}e_{N+1}(x)=-\dfrac{f^{\prime}(a)\varphi^{(N+1)}(0)}{(N+1)!}.$ (12)
    From this, if $\varphi^{(N+1)}(0)=0$ holds, then the statement is proven.
    Let $k<N+1$ be a positive integer and observe the limit $\lim_{x\rightarrow a}e_{N+1}$ after $k$-many applications of the l’Hospital rule. Note that apart from the terms containing $f^{(N+1)}(a)$, the numerator is the same as the numerator of $e_{N}$ after the same number of applications of the rule. This means that by the assumption of the induction step, these terms tend to zero. Moreover, the terms containing $f^{(N+1)}(a)$ initially had the ($N+1$)th power of $\varphi(x-a)$, meaning that after $k$-many ($k<N+1$) differentiations every term will contain at least one $\varphi(x-a)$ term, meaning that this also tends to zero (since $\varphi(0)=0$). The same can also be applied to the denominator, meaning that (at least) one additional l’Hospital rule can be applied. This means that in order to calculate $\lim_{x\rightarrow a}e_{N+1}$, $N+1$-many l’Hospital rules have to be applied.
    Now, observe the form of the limit after ($N+1$)-many differentiations. In the numerator, almost every term has at least either one $\varphi(x-a)$ term or one $\varphi^{(\ell)}(x-a)$ ($2\leq\ell\leq N$) term, thus, by the assumptions of (C1), all of these tend to zero. The three exceptions are:
    • the term $f^{(N+1)}(x)$,
    • the term $-f^{\prime}(a)\varphi^{(N+1)}(x-a)$,
    • the term $\left(\varphi^{\prime}(x-a)\right)^{N+1}$. (The case $\left(\varphi^{\prime}(x-a)\right)^{\ell}$ where $\ell<N+1$ is not possible since we performed $N+1$-many differentiations, so the sum of the orders of derivatives in a product should also be $N+1$.) Such a term is multiplied by $-f^{(N+1)}(a)$ (it cannot appear elsewhere since the number of terms in a product remains the same during differentiation).
    Note that the denominator is similar to the terms in the numerator multiplied by $f^{(N+1)}(a)$: in the denominator, instead of $f^{(N+1)}(a)$, they are multiplied by $(N+1)!$. Consequently, a similar argument can also be made there: the only nonzero term is $\left(\varphi^{\prime}(x-a)\right)^{N+1}$.
    In conclusion,
    $\lim_{x\rightarrow a}e_{N+1}(x)=\dfrac{f^{(N+1)}(a)-f^{\prime}(a)\varphi^{(N+1)}(0)-f^{(N+1)}(a)\left(\varphi^{\prime}(0)\right)^{N+1}}{(N+1)!\left(\varphi^{\prime}(0)\right)^{N+1}}=-\dfrac{f^{\prime}(a)\varphi^{(N+1)}(0)}{(N+1)!}.$
    Therefore, claim (12) is proven.
    Now we show that $e_{n}(x)=O(|x-a|)$. For this, it is enough to prove that $\dfrac{|e_{n}(x)|}{|x-a|}$ is bounded if $|x-a|$ is smaller than a given constant. Let us observe that
    $\dfrac{|e_{n}(x)|}{|x-a|}=\dfrac{\left|f(x)-f(a)-f^{\prime}(a)\varphi(x-a)-\dfrac{f^{\prime\prime}(a)}{2}\left(\varphi(x-a)\right)^{2}-\dots\dfrac{f^{(n)}(a)}{n!}\left(\varphi(x-a)\right)^{n}\right|}{|x-a|}$ (13)
    By condition (C1), we get that the ’standard’ Taylor series of $\varphi(x-a)$ around 00 has the form $\varphi(x-a)=(x-a)+O(|x-a|^{n+1})$, meaning that the numerator of (13) has the form
    $\left|f(x)-f(a)-f^{\prime}(a)\left((x-a)+O(|x-a|^{n+1})\right)-\dfrac{f^{\prime\prime}(a)}{2}\left((x-a)+O(|x-a|^{n+1})\right)^{2}-\dots\right.$
    $\left.\dots-\dfrac{f^{(n)}(a)}{n!}\left((x-a)+O(|x-a|^{n+1})\right)^{n}\right|=$
    $=\left|\dfrac{f^{(n+1)}(\xi_{1})}{(n+1)!}|x-a|^{n+1}-f^{\prime}(a)\dfrac{\varphi^{(n+1)}(\xi_{2})}{(n+1)!}|x-a|^{n+1}+O(|x-a|^{n+2})\right|,$
    where $\xi_{1},\xi_{2}\in(a,x)$ and we used the Lagrange forms of the remainders. Then, by dividing this term by $|x-a|$, we get an expression that is bounded if $|x-a|$ is small enough.
@@
    Here we show the (C2$\Rightarrow$(C1) direction.
    We again proceed by induction. For $n=0$, we have no condition.
    For $n=1$, we assume proceeding towards contradiction that $\varphi(0)\neq 0$. Then, (8) has the form
    $f(x)=f(a)+f^{\prime}(a)\varphi(x-a)+e_{1}(x)\varphi(x-a).$
    By taking the limit $x\rightarrow a$ on both sides, we get a contradiction.
    Now, assume that $\varphi^{\prime}(0)\neq 1$. Similarly to the proof of the first part of the statement, we get that
    $0=\lim_{x\rightarrow a}e_{1}(x)=\dfrac{f^{\prime}(a)-f^{\prime}(a)\varphi^{\prime}(0)}{\varphi^{\prime}(0)}=f^{\prime}(a)\left(\dfrac{1}{\varphi^{\prime}(0)}-1\right).$
    This can only hold for any arbitrary $f$ (which has the properties specified in the statement of the lemma) if $\varphi^{\prime}(0)=1$.
    In order to prove the induction step, let us assume that the statement is true for $n=N$. As in the previous half of the proof, one can show that
    $\lim_{x\rightarrow a}e_{N+1}(x)=-\dfrac{f^{\prime}(a)\varphi^{(N+1)}(0)}{(N+1)!},$
    which again can only be zero if $\varphi^{(N+1)}(0)=0$. ∎
Remark 5.
Claim (12) can also be proven by applying the Faà di Bruno formula (see e.g. [54]) to $g(h(x))$ with $g(x)=x^{N+1}$ and $h(x)=\varphi(x-a)$. Moreover, the direction (C1) $\Rightarrow$ (C2) is also the consequence of the proof of the $e_{n}(x-a)=O(|x-a|)$ claim, but the longer version is needed for the (C2) $\Rightarrow$ (C1) direction.
@@
    Let us observe the following linear difference operator $L$ (see e.g. [51]):
@@
    $=y(t)\sum_{j=0}^{k}\alpha_{j}+\sum_{q=1}^{p}\dfrac{y^{(q)}(t)}{q!}\sum_{j=0}^{k}\alpha_{j}\left(\varphi(j\Delta t)\right)^{q}-\varphi(\Delta t)\sum_{r=0}^{p}\dfrac{y^{(r+1)}(t)}{r!}\sum_{j=0}^{k}\beta_{j}\left(\varphi(j\Delta t)\right)^{r}+O((\Delta t)^{p+1})=$
    $=\sum_{q=1}^{p}\dfrac{y^{(q)}(t)}{q!}\sum_{j=0}^{k}\alpha_{j}\left(\varphi(j\Delta t)\right)^{q}-\varphi(\Delta t)\sum_{r=1}^{p+1}r\dfrac{y^{(r)}(t)}{r!}\sum_{j=0}^{k}\beta_{j}\left(\varphi(j\Delta t)\right)^{r-1}+O((\Delta t)^{p+1}),$
@@
    $=\sum_{q=1}^{p}\dfrac{y^{(q)}(t)}{q!}\left(\Delta t\right)^{q}\left(\sum_{j=0}^{k}\alpha_{j}j^{q}-q\sum_{j=0}^{k}\beta_{j}j^{q-1}\right)+O((\Delta t)^{p+1}).$
@@
    Let us apply the ’standard’ Taylor expansion for the functions $y(t+j\Delta t)$ and $y^{\prime}(t+j\Delta t)$ at points $x=t+j\Delta t$:
@@
    $=\sum_{j=0}^{k}\alpha_{j}y(t)+y^{\prime}(t)\Delta t\sum_{j=0}^{k}\left(j\alpha_{j}-\varphi^{\prime}(0)\beta_{j}\right)+O((\Delta t)^{2})=$
    $=y^{\prime}(t)\Delta t\sum_{j=0}^{k}\left(j\alpha_{j}-\varphi^{\prime}(0)\beta_{j}\right)+O((\Delta t)^{2}),$
@@
    As before, we apply the ’standard’ Taylor expansion to functions $y(t+j\Delta t)$ and $y^{\prime}(t+j\Delta t)$ at points $x=t+j\Delta t$, resulting in
    $L(y,t,\Delta t)$ $=\sum_{j=0}^{k}\left(\alpha_{j}\sum_{q=0}^{N}\dfrac{\left(j\Delta t\right)^{q}}{q!}y^{(q)}(t)-\right.$
@@
    $=\sum_{q=1}^{N}\dfrac{y^{(q)}(t)}{q!}\left(\Delta t\right)^{q}\left(\sum_{j=0}^{k}\alpha_{j}j^{q}-q\sum_{j=0}^{k}\beta_{j}j^{q-1}\right)-\dfrac{\varphi^{(N)}(0)}{N!}(\Delta t)^{N}\sum_{j=0}^{k}\beta_{j}y^{\prime}(t)+O((\Delta t)^{N+1})=$
    $=-\dfrac{\varphi^{(N)}(0)}{N!}(\Delta t)^{N}\sum_{j=0}^{k}\beta_{j}y^{\prime}(t)+O((\Delta t)^{N+1}),$ (14)
@@
Let us consider the numerical method (10) and observe its part concerning the element $u_{k}$:
@@
We prove the statement by induction: let us assume that $u_{k}^{n+1}\geq\min_{\ell=n-s+1,\;n-s+2,\;\dots,\;n}\{u_{k}^{\ell}\}$ for $n=s-1,\;s,\;\dots,\;N-1$ and our goal is to show that $u_{k}^{N+1}\geq\min_{\ell=N-s+1,\;N-s+2,\;\dots,\;N}\{u_{k}^{\ell}\}$ ($N+1\leq\mathcal{N}$).
We again consider the numerical method (10) and observe its part concerning the element $u_{k}$ (see (15) in the proof of Theorem 3). Since the monotone increase holds for the forward Euler method if $\Delta t_{FE}\leq B_{FE}$, for $\Delta t_{FE}\leq B_{FE}$ we have
@@
$u_{k}^{N+1}=\sum_{j=1}^{s}\tilde{\alpha}_{j}\left(u_{k}^{N+1-j}+\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}f_{k}(u^{N+1-j})\right)\geq\sum_{j=1}^{s}\tilde{\alpha}_{j}u_{k}^{N+1-j}.\geq$
@@
We prove the statement by induction: let us assume that $P(u_{1}^{n},\;u_{2}^{n},\;\dots,u_{m}^{n})=\mathcal{M}$ for $n=s,\;s+1,\;\dots,\;N$ and our goal is to show that $P(u_{1}^{N+1},\;u_{2}^{N+1},\;\dots,u_{m}^{N+1})=\mathcal{M}$ ($N+1\leq\mathcal{N}$).
@@
$=\sum_{j=1}^{s}\tilde{\alpha}_{j}\sum_{\ell=1}^{m}\gamma_{\ell}\left(u_{\ell}^{N+1-j}+\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}f_{\ell}(u^{N+1-j})\right)=\sum_{j=1}^{s}\tilde{\alpha}_{j}M=M,$

### v2

@@ 3.3 The choice of function $\varphi$
We should also note here that in [51], Dang and Hoang demonstrated a different construction of functions fulfilling these conditions up to an arbitrary order.
Remark 9.
It should be noted that although the condition $\varphi(x)\leq\mathcal{B}$ ensures the preservation of the aforementioned qualitative properties, such choices where the function $\varphi(x)$ is close to the constant function $\varphi_{0}(x)\equiv\mathcal{B}$ might still result in considerably large errors when using large timesteps, similarly as in the case of “standard” methods. The benefit of these nonstandard methods is that even for larger timesteps, the numerical solution behaves in a qualitatively reasonable way, although the errors might still be relatively big. For smaller timesteps, the methods behave as expected resulting in small errors, like other, high order standard methods.
In Section 5 we perform several numerical experiments comparing these choices. There, it turns out that it is usually beneficial to apply such $\varphi$ functions which allow a higher order of convergence (e.g. $\varphi_{8}$), since their application generally results in smaller errors, and sometimes even higher rates of convergence than the one which is guaranteed by the multistep method itself, see the third order methods in Section 5.2 that attain an order of four. Numerical tests (see Table 1) indicate that there is no significant difference between the runtime of methods using e.g. $\varphi_{3}$ or $\varphi_{8}$, so it is advised to use the latter: the functions that require more time to compute are $\varphi_{1}$ and $\varphi_{2}$, but their application is not recommended since they decrease the order of the methods significantly.
$\varphi_{1}$ $\varphi_{2}$ $\varphi_{3}$ $\varphi_{4}$ $\varphi_{5}$ $\varphi_{6}$ $\varphi_{7}$ $\varphi_{8}$
runtime (s) $7.2136$ $7.0121$ $0.4383$ $0.4377$ $0.4358$ $0.4447$ $0.4382$ $0.4410$
Table 1: The time (seconds) it takes to calculate the value of a given function $10^{9}$-many times with $\mathcal{B}=0.1$ and $x=0.1$. The tests were made by using Matlab R2024b.
@@ 4 Proofs of the statements of Section 3
The main idea of the proof is that condition (C1) is equivalent to the following condition:
  1. (C3)
    the (standard) Taylor series of function $\varphi$ around 0 at point xax-a is in the form
    $\varphi(x-a)=(x-a)+O((x-a)^{n+1}).$
It is easy to see that if (C1) holds, then by the standard Taylor expansion of $\varphi$ around 0 we get (C3). Moreover, (C3) can only hold if (C1) is true (again, by the form of the standard Taylor expansion of $\varphi$).
We prove the two directions of the lemma separately.
@@
[…]
    $f(x)=\sum_{k=0}^{n}\dfrac{f^{(k)}(a)}{k!}(x-a)^{k}+O((x-a)^{n+1}).$ (12)
    Next, we substitute the statement of (C3) into the right-hand side of (C2):
    $\sum_{k=0}^{n}\dfrac{f^{(k)}(a)}{k!}\left(\varphi(x-a)\right)^{k}+O(x-a)\left(\varphi(x-a)\right)^{n}=$
    $\sum_{k=0}^{n}\dfrac{f^{(k)}(a)}{k!}\left((x-a)+O((x-a)^{n+1})\right)^{k}+O(x-a)\left((x-a)+O((x-a)^{n+1})\right)^{n}=$
    $\sum_{k=0}^{n}\dfrac{f^{(k)}(a)}{k!}\left(x-a)\right)^{k}+O\left((x-a)^{n+1}\right)=f(x),$
    where we used (12).
@@
[…]
    $f(x)=\sum_{k=0}^{n}\dfrac{f^{(k)}(a)}{k!}\left(\varphi(0)+\varphi^{\prime}(0)(x-a)+\sum_{\ell=2}^{n}\dfrac{\varphi^{(\ell)}(0)}{\ell!}(x-a)^{\ell}+O((x-a)^{n+1})\right)^{k}+$
    $+O(x-a)\left(\varphi(0)+\varphi^{\prime}(0)(x-a)+\sum_{\ell=2}^{n}\dfrac{\varphi^{(\ell)}(0)}{\ell!}(x-a)^{\ell}+O((x-a)^{n+1})\right)^{n}.$
    By comparing this expression with (12), they can only hold if
    $\left(\varphi(0)+\varphi^{\prime}(0)(x-a)+\sum_{\ell=2}^{n}\dfrac{\varphi^{(\ell)}(0)}{\ell!}(x-a)^{\ell}\right)^{k}=(x-a)^{k},\qquad k=1,\;2,\;\dots\;n,$
    which can only hold if (C1) is true.∎
@@
    Let us consider the following linear difference operator $L$ (see e.g. [53]):
@@
    $y(t)\sum_{j=0}^{k}\alpha_{j}+\sum_{q=1}^{p}\dfrac{y^{(q)}(t)}{q!}\sum_{j=0}^{k}\alpha_{j}\left(\varphi(j\Delta t)\right)^{q}-\varphi(\Delta t)\sum_{r=0}^{p}\dfrac{y^{(r+1)}(t)}{r!}\sum_{j=0}^{k}\beta_{j}\left(\varphi(j\Delta t)\right)^{r}+O((\Delta t)^{p+1})=$
    $\sum_{q=1}^{p}\dfrac{y^{(q)}(t)}{q!}\sum_{j=0}^{k}\alpha_{j}\left(\varphi(j\Delta t)\right)^{q}-\varphi(\Delta t)\sum_{r=1}^{p+1}r\dfrac{y^{(r)}(t)}{r!}\sum_{j=0}^{k}\beta_{j}\left(\varphi(j\Delta t)\right)^{r-1}+O((\Delta t)^{p+1}),$
@@
    $\sum_{q=1}^{p}\dfrac{y^{(q)}(t)}{q!}\left(\Delta t\right)^{q}\left(\sum_{j=0}^{k}\alpha_{j}j^{q}-q\sum_{j=0}^{k}\beta_{j}j^{q-1}\right)+O((\Delta t)^{p+1}).$
@@
    Let us apply the “standard” Taylor expansion for the functions $y(t+j\Delta t)$ and $y^{\prime}(t+j\Delta t)$ at points $x=t+j\Delta t$:
@@
    $\sum_{j=0}^{k}\alpha_{j}y(t)+y^{\prime}(t)\Delta t\sum_{j=0}^{k}\left(j\alpha_{j}-\varphi^{\prime}(0)\beta_{j}\right)+O((\Delta t)^{2})=$
    $y^{\prime}(t)\Delta t\sum_{j=0}^{k}\left(j\alpha_{j}-\varphi^{\prime}(0)\beta_{j}\right)+O((\Delta t)^{2}),$
@@
    As before, we apply the “standard” Taylor expansion to functions $y(t+j\Delta t)$ and $y^{\prime}(t+j\Delta t)$ at points $x=t+j\Delta t$, resulting in
[…]
@@
    $\sum_{q=1}^{N}\dfrac{y^{(q)}(t)}{q!}\left(\Delta t\right)^{q}\left(\sum_{j=0}^{k}\alpha_{j}j^{q}-q\sum_{j=0}^{k}\beta_{j}j^{q-1}\right)-\dfrac{\varphi^{(N)}(0)}{N!}(\Delta t)^{N}\sum_{j=0}^{k}\beta_{j}y^{\prime}(t)+O((\Delta t)^{N+1})=$
    $\dfrac{\varphi^{(N)}(0)}{N!}(\Delta t)^{N}\sum_{j=0}^{k}\beta_{j}y^{\prime}(t)+O((\Delta t)^{N+1}),$ (13)
@@
Let us consider the numerical method (10) and analyze its part concerning the element $u_{k}$:
@@
We prove the statement by induction: let us assume that
$u_{k}^{n+1}\geq\min_{\ell=n-s+1,\;n-s+2,\;\dots,\;n}\{u_{k}^{\ell}\}\qquad\text{for}\;n=s-1,\;s,\;\dots,\;N-1$
and our goal is to show that $u_{k}^{N+1}\geq\min_{\ell=N-s+1,\;N-s+2,\;\dots,\;N}\{u_{k}^{\ell}\}$ ($N+1\leq\mathcal{N}$).
We again consider the numerical method (10) and analyze its part concerning the element $u_{k}$ (see (14) in the proof of Theorem 3). Since the monotone increase holds for the forward Euler method if $\Delta t_{FE}\leq B_{FE}$, for $\Delta t_{FE}\leq B_{FE}$ we have
@@
$u_{k}^{N+1}=\sum_{j=1}^{s}\tilde{\alpha}_{j}\left(u_{k}^{N+1-j}+\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}f_{k}(u^{N+1-j})\right)\geq\sum_{j=1}^{s}\tilde{\alpha}_{j}u_{k}^{N+1-j}\geq$
@@
We prove the statement by induction: let us assume that $P(u_{1}^{n},\;u_{2}^{n},\;\dots,u_{m}^{n})=\mathcal{M}$ for $n=s,\;s+1,\;\dots,\;N$ and our goal is to show that $P(u_{1}^{N+1},\;u_{2}^{N+1},\;\dots,u_{m}^{N+1})=\mathcal{M}$ ( here $N+1\leq\mathcal{N}$).
@@
$\sum_{j=1}^{s}\tilde{\alpha}_{j}\sum_{\ell=1}^{m}\gamma_{\ell}\left(u_{\ell}^{N+1-j}+\varphi(\Delta t)\dfrac{\tilde{\beta}_{j}}{\tilde{\alpha}_{j}}f_{\ell}(u^{N+1-j})\right)=\sum_{j=1}^{s}\tilde{\alpha}_{j}M=M,$